OpenStudy (anonymous):

3. Mark paid a $40 one-time fee to join his local gym and his membership is $15 a month. Write an expression to find the total amount at the gym after n months. A. 15n + 40 B. 40(n + 15) C. 40n D. 40 + 15n

hartnn (hartnn):

15$ a month means 15n for 'n' months ok? one time fees will be added to ths so 40+15n

hartnn (hartnn):

and whats the difference between option A and option D ??

OpenStudy (anonymous):

i dont know exactly .other than there inversed i think is the word?

OpenStudy (anonymous):

and that was my option! :)

hartnn (hartnn):

they both are exactly the same 15n+40=40+15n = correct answer

OpenStudy (anonymous):

so i was right then D?

hartnn (hartnn):

i already said BOTH A and D are correct! and exactly same

hartnn (hartnn):

so, did u get it right ?

OpenStudy (anonymous):

im still working on a crap ton more questions ill definitely let you know when i get done.lol

hartnn (hartnn):

okay :) lets go to next question ^_^